OpenStudy (anonymous):

5*5=25 9*10=90 so answer is 9025

OpenStudy (anonymous):

Is this 95 times 95?

OpenStudy (anonymous):

s

OpenStudy (anonymous):

there is something worth remembering when multiplying numbers like 3.3 * 3.5 and 95 * 95: 3.5 * 3.5 = 3 * 4 + 0.25 = 12.25 6.5 * 6.5 = 6*7 + 0.25 = 42.25 95* 95 = 90 * 100 + 25 = 9025 105 * 105 = 100*110 + 25 = 11025